### v1

@@ 2.5 Emergent phenomenon
When building domain-specific foundation models, developers must select an appropriate model scale, considering both the requirements of potential downstream tasks and user prompting patterns. An increase in the number of model parameters raises the demand for computational resources and the risk of overfitting. Consequently, there is a limit to the extent that parameters can be increased. The trade-offs highlighted by this phenomenon are crucial for model deployment, offering significant insights for both model design and application.
@@ 3 Key technologies for building domain-specific foundation models
This section delves into the technical pathways toward customizing domain-specific foundation models. We will provide a detailed explanation of how to flexibly select and combine the appropriate modules from five key components — modality encoders, input projectors, backbone calculators, output projectors, and modality decoders based on the specific requirements of different domains. Additionally, we will analyze concrete cases to help readers better understand and apply the methodologies discussed in this section.
We can categorize the customization of domain-specific foundation models into three levels, ranging from low to high level of customization (i.e., from high to low reliance on general-purpose foundation models or pre-trained modules):
@@
    Domain-specific enhancement based on general-purpose foundation models.
@@
    Customization of the foundation model based on pre-trained modules.
@@
    Construction of the foundation model without pre-trained modules.
Table 3 summarizes the characteristics of these three domain-specific foundation model customization methods.
Table 3: Customization methods of domain-specific foundation models
@@
Based on general-purpose foundation models Low, only the input method of domain knowledge is customized Low Low Low
@@
3.1 Domain-specific enhancement based on general-purpose foundation models
General-purpose foundation models offer comprehensive capabilities, making them suitable for a wide array of task scenarios. When such a general-purpose foundation model can fully handle the required data modalities, it renders unnecessary any modifications to its underlying architecture for model developers. Instead, they can focus on implementing domain-specific enhancements.
Depending on whether the domain-specific enhancement requires altering the parameters of the foundation model, we can further divide this into two types: plug-and-play domain-specific enhancement and fine-tuning-based domain-specific enhancement. Table 4 categorizes and summarizes the methods of domain-specific enhancement based on a full-architecture general-purpose foundation model.
Table 4: Specific-domain enhancement with the entire general-purpose foundation models
@@ 3.1.1 Plug-and-play domain-specific enhancement
The generality, generalization capabilities, and reasoning abilities of general-purpose foundation models enable them to serve as the foundation for domain-specific models. To achieve plug-and-play domain enhancement without modifying the parameters of the foundation model, two approaches can be utilized: leveraging existing knowledge or embedding new knowledge. The first approach aims to leverage domain knowledge already stored within the general-purpose foundation model, as illustrated in Figure 5 (a). The second approach, embedding new knowledge, involves equipping the foundation model with the ability to handle domain tasks by introducing domain-specific knowledge. This can be further divided into two methods: embedding knowledge through prompts and embedding knowledge via an external knowledge base. These methods are depicted in Figure 5 (b) and (c) respectively. The following sections will provide a detailed explanation of these techniques.
@@
    Invoking existing knowledge for domain enhancement: During the training process, a general-purpose foundation model may have already enclosed domain knowledge. Prompt tuning improves prompts to better invoke the model’s inherent domain knowledge, where ”tuning” refers to the optimization of prompts. Specifically, it involves inserting a carefully crafted prompt into the input data as context to improve the generated output. These carefully crafted prompts can be natural language descriptions, examples, rules, or other text or embedding vectors that guide the model to understand the task requirements. When generating outputs, the model will consider these carefully crafted prompts to produce task-relevant results. Prompt tuning can be categorized into hard prompts and soft prompts:
@@
      Hard prompts: Hard prompt methods are common techniques in natural language processing (NLP). They guide the language model’s output using interpretable and reusable handcrafted words and tokens. Hard prompts are typically manually designed and tailored for specific tasks, making them difficult to modify. PET (Pattern Exploiting Training) [62] is a classic hard prompt learning method that models questions as cloze tasks and optimizes the final output words. This method trains the model on a small amount of supervised data and performs ensembling predictions on unsupervised data to guide the model.
@@
      Soft prompts: Designing hard prompts requires experimental exploration and expertise, and manually designed prompts may not align well with model’s data processing methods. To simplify this process and enhance the flexibility of prompt tuning, researchers proposed soft prompt-based tuning methods. Prefix Tuning [63] is a form of soft prompt tuning that adapts to specific downstream tasks by adding learnable prefix vectors (soft prompts) to the beginning of the input sequence. These prefix vectors, as part of the input, guide the model’s output to meet task requirements. The advantage of prefix tuning is that it only updates these prefix vectors rather than the model’s parameters, significantly reducing computational and storage resource demands while retaining the rich knowledge learned by the pre-trained model. Building on prefix tuning, researchers introduced the P-tuning method [64]. P-tuning replaces fixed or manually designed words and tokens with learnable soft prompts. Its core idea is to treat prompts as part of the model that can be learned, allowing the model to learn not only how to respond to given tasks but also how to generate the best prompts. These soft prompts are typically a series of embedding vectors processed alongside the actual text input. Through end-to-end training, the model automatically learns to adjust these embedding vectors for better task performance. P-tuning combines the parameter efficiency of prefix tuning with the flexibility of traditional hard prompt tuning. Soft prompts give the model greater freedom to generate answers, potentially producing more diverse outputs but also increasing the risk of generating inaccurate or irrelevant responses.
@@
    Knowledge embedding for domain enhancement: When the existing knowledge of a general-purpose foundation model is insufficient to solve domain tasks, introducing new knowledge by embedding additional background information can achieve higher quality output. This method is known as knowledge embedding for domain enhancement.
@@
      • Resource consumption of the attention mechanism: The attention mechanism is the core of Transformer models, allowing the model to compute attention weights for each element in the sequence. However, as the sequence length increases, the computational complexity and memory requirements of this mechanism grow quadratically, leading to significant resource consumption.
      • Long-distance dependency issue: When handling long sequences, the Transformer needs to span a large number of input tokens, which often results in problems such as gradient vanishing or exploding. This makes it difficult for the model to capture dependencies between elements that are far apart in the sequence.
      To address the above issues, Lossless Long Text technology has emerged. It aims to enhance the model’s ability to process long texts that exceed its input length limitations, allowing users to input a large amount of domain knowledge directly through prompts into the large language model as contextual information for domain enhancement. Lossless Long Text technology expands the long text input capability of large language models in two directions: extrapolation and interpolation:
@@
      Knowledge embedding by external knowledge base: In practical applications, users may not be able to provide sufficient domain knowledge to enhance a general-purpose foundation model. To address this issue, model deployers can augment the general-purpose foundation model with a dedicated domain knowledge base. This method allows the general-purpose foundation model to reference this external knowledge base when generating answers or performing tasks, thereby obtaining the necessary domain information and context to provide more accurate and targeted responses or solutions. Retrieval-Augmented Generation (RAG) [67, 1, 81] technology was developed to achieve this purpose. RAG technology aims to enhance the language model’s generation capabilities by leveraging an external document base without retraining the model. It is particularly suitable for tasks requiring a customizable dynamic knowledge base, such as question answering, text summarization, and fact-checking. The core of RAG technology is the integration of a retrieval component that can quickly find information relevant to the current task in a large document database during the generation process. Once the relevant documents are retrieved, this information is used as additional contextual information to aid the generation process. The advantage of RAG technology is that it combines the generation capabilities of large language models with the knowledge provided by external retrieval systems, without requiring domain knowledge themselves. Furthermore, because the external knowledge base can be replaced as needed, RAG technology offers high flexibility and adaptability.
@@
      In RAG, the two core challenges are how to retrieve useful information from the database and how to organize that information to augment the generation of foundation models. For the retrieval task, the earliest and most naive approach uses sparse retrieval that directly matches data based on raw data like BM25 [82]. Inspired by the field of information retrieval, dense retrieval like DPR [83] has been proposed, which projects raw data into a high-dimensional space, potentially helping to capture semantic similarity better. However, a significant challenge is that even if the retrieved data is highly related to the original input in the semantic space, we cannot guarantee that it will help the model generation, due to issues like polysemy. To address this problem, some researchers have introduced other technologies like knowledge graphs to aid the retrieval process [68][84] showed that adding noise can help enhance the model performance compared to traditional dense retrieval methods. Additionally, the organization of the retrieved information can also directly influence the output quality. For example, providing a ranking of the retrieved data [69] can improve model performance.
@@
While the aforementioned technologies were initially proposed to enhance large language models in specific domains, their applications are not limited to language models. With the development of foundation model domains, these technologies are expected to be extended to foundation models of other modalities.
@@ 3.1.2 Domain-specific enhancement based on fine-tuning
When plug-and-play domain enhancement techniques are difficult to implement or require embedding too much domain knowledge into the general-purpose foundation model, or when deep modifications to the general-purpose foundation model are necessary, we can turn to the strategy of domain enhancement based on fine-tuning. This strategy aims to customize the required domain-specific foundation model while preserving the pre-trained knowledge of the general-purpose foundation model as much as possible by specific domain enhancement [85].
Fine-tuning techniques can be divided into three main types: adapter-based fine-tuning, low-rank matrix decomposition-based fine-tuning, and full-parameter fine-tuning. Figure 6 (a), (b) and (c) illustrate these three technical pathways respectively. In following sections, this paper will elaborate on these techniques, sorted from low to high resource requirements and complexity needed for fine-tuning.
@@
    Adapter-based fine-tuning: Adapter-based fine-tuning [70] is a method that inserts small trainable adapter modules into pre-trained models, aiming to efficiently adapt the model to specific downstream tasks. During fine-tuning, only parameters of adapter modules are updated, while the original parameters of the pre-trained model remain unchanged, reducing computational resources and storage requirements while retaining the rich knowledge learned by the model during pre-training. AdapterFusion [71] is an extension of adapter-based fine-tuning that allows the model to learn multiple tasks or adapt to various data distributions simultaneously by fusing multiple adapter modules, with each adapter module focusing on capturing task-specific features. Infused Adapter by Inhibiting and Amplifying Inner Activations (IA3) [72] scales the activation layers by injecting learned vectors into the attention and feed-forward modules of the Transformer architecture. Since these learned vectors are only trainable parameters during fine-tuning, IA3 significantly reduces the number of trainable parameters compared to traditional adapter-based fine-tuning, thereby reducing training costs and improving training efficiency. Additionally, IA3 does not introduce inference latency because its adapter weights can be merged with the foundation model while maintaining the flexibility and adaptability of the model, allowing customized fine-tuning for different tasks and datasets. [73] proposes a method called Model Reprogramming that converts both the input and output domains of a foundation model by adding two adapter layers, enabling more flexible adaptation functionality.
@@
    Low-rank matrix decomposition based fine-tuning: Fine-tuning based on low-rank matrix decomposition reduces the number of parameters that need to be updated by decomposing the weight matrices in the pre-trained model into the product of low-rank matrices. Low-rank matrix decomposition can capture the most important information in the weight matrices while keeping the original pre-trained parameters unchanged during fine-tuning, updating only low-rank decomposition matrices, thus reducing the computational and storage requirements during fine-tuning. This method improves fine-tuning efficiency while maintaining or approaching the performance of full-parameter fine-tuning.
@@
      Low-Rank Adaptation (LoRA): Low-Rank Adaptation (LoRA) [74] achieves efficient fine-tuning performance by decomposing model parameters into the product of low-rank matrices using singular value decomposition (SVD). The principle of LoRA can be represented by the following equation:
@@
      where $W_{\text{old}}$ is the original weight matrix, and $\Delta W$ is the low-rank update matrix, which can be constructed by selecting the singular vectors corresponding to smaller singular values of $W_{\text{old}}$. The singular value decomposition of $W_{\text{old}}$ is given by:
      $\Delta W=U\Sigma V^{T}\text{.}$ (11)
      Here, $U$ and $V$ are matrices obtained from the SVD of $W_{\text{old}}$, and $\Sigma$ is a diagonal matrix containing the most important singular values of $W_{\text{old}}$. By updating only the parameters in $U$, $\Sigma$, and $V$, LoRA efficiently fine-tunes the model.
@@
      Researchers have also found that LoRA’s continuous pre-training performance is unsatisfactory on some large-scale datasets. Thus, Layerwise Importance Sampled AdamW (LISA) [87] strategy is proposed, where the weight norm distribution of different layers exhibits uncommon skewness. LISA adopts importance sampling strategy by randomly activating different layers in the foundation model for optimization. Specifically, LISA consistently updates the bottom embedding layers and the top linear head while randomly updating a small number of intermediate self-attention layers. This method, with memory consumption comparable to LoRA, outperforms LoRA and even full-parameter fine-tuning in various downstream fine-tuning tasks.
@@
      $W_{\text{new}}=W_{\text{old}}\odot\Delta W\text{,}$ (12)
      where $W_{\text{old}}$ is the original weight matrix, and $\Delta W$ is the update matrix approximated by a low-rank matrix. The update matrix $\Delta W$ can be further decomposed into the Hadamard product of two low-rank matrices:
      $\Delta W=L_{1}\odot L_{2}\text{.}$ (13)
      Here, $L_{1}$ and $L_{2}$ are two low-rank matrices that adjust elements of the original weight matrix by learning key information extracted from the input data. By updating only parameters in $L_{1}$ and $L_{2}$, LoHa achieves efficient fine-tuning of the model while maintaining its adaptability to new tasks.
@@
      Low-Rank Kronecker Product (LoKr): Low-Rank Kronecker Product (LoKr) [76] is another parameter-efficient fine-tuning method that emerged after LoHa. LoKr utilizes the properties of the Kronecker product to expand the dimensions of the weight matrix while keeping the increase in the number of parameters within a manageable range. The Kronecker product allows the model to learn complex interactions across different dimensions, which is particularly useful for capturing high-order relationships in the input data. The updating process of LoKr can be represented as:

### v2

@@ 2.3 Gain of scaling up
When building domain-specific FMs, developers must select an appropriate model scale, considering both the requirements of potential downstream tasks and user prompting patterns. An increase in the number of model parameters raises the demand for computational resources and the risk of overfitting. Consequently, there is a limit to the extent that parameters can be increased. The trade-offs highlighted by this phenomenon are crucial for model deployment, offering significant insights for both model design and application.
2.4 Performance comparison
In this section, we present the evaluation results of current multi-modality FMs to assist readers in selecting the appropriate model for their needs. We summarize some common vision-language tasks in Table 3 according to  [26, 61, 62, 63]. We select three mainstream vision-language tasks, including:
  • Image Captioning: This task involves generating descriptive captions for images, demonstrating models’ ability to understand visual content.
  • Image Question Answering: In this task, models are required to answer questions about the content of images, demonstrating their ability to comprehend and reason about visual information.
  • Benchmark Toolkit: This refers to a collection of standardized metrics and datasets used to evaluate the performance of multi-modal models across various tasks like conversation, complex reasoning, and detailed description, ensuring a consistent assessment framework.
In summary, InternVL-Chat [63] has the best results in Image Captioning and Image Question Answering tasks. While considering language tasks such as Benchmark Toolkits, the best results are provided by NExT-GPT [26].
Table 3: Performance of various large multi-modality models on vision-language tasks
Model LLM Image Captioning Image Question Answering Benchmark Toolkit NoCaps [64] Flickr30K [65] COCO [66] VQAv2 [67] GQA [68] VizWiz [69] SQAI [70] VQAT [71] OKVQA [72] POPE [73] MME [31] MMB [74] LLAVAW [23] SEED [75] MM-Vet [76] QBench [77] mPLUG-Owl  [78] LLaMA-7B  [10] 117.0 80.3 119.3 - - 39.0 - - - - - 46.6 - 34.0 - - Emu  [79] LLaMA-7B  [10] - - 117.7 40.0 - 35.4 - - 34.7 - - - - - - - I-80B  [80] LLaMA-65B  [10] 65.0 53.7 91.8 60.0 45.2 36.0 - 30.9 - - - 54.5 - - - - LLAVA  [23] LLaMA2-7B  [11] 120.7 82.7 - - - - - - - - - 36.2 - - - - MobileVLM  [81] MobileLLaMa-2.7B  [82] - - - - 59.0 - 61.0 47.5 - 84.9 1288.9 59.6 - - - - DREAMLLM  [83] Vicuna-7B  [84] - - 115.4 56.6 - 45.8 - - 44.3 - - 49.9 - - - - Video-LLAVA  [85] Vicuna-7B  [84] - - - 74.7 - 48.1 - - - - - 60.9 - - - - NExT-GPT  [26] Vicuna-7B  [84] 123.7 84.5 124.9 66.7 - 48.4 - - 52.1 - - 58.0 - 57.5 - - ShareGPT4V-7B  [61] Vicuna-7B  [84] - - - 80.6 - 57.2 68.4 - - - 1567.4 68.8 72.6 69.7 37.6 63.4 LLAVA-1.5  [86] Vicuna-7B  [84] - - - 78.5 62.0 50.0 - 58.2 - 85.9 1510.7 - - - - - LLAVA-1.5  [86] Vicuna-13B  [84] - - - 80.0 63.3 53.6 71.6 61.3 - 85.9 1531.3 67.7 70.7 68.2 35.4 62.1 InstructBLIP  [87] Vicuna-7B  [84] 123.1 82.4 102.2 - - 34.5 60.5 50.1 33.9 - - 36.0 60.9 53.4 26.2 56.7 InstructBLIP  [87] Vicuna-13B  [84] 121.9 82.8 - - 49.5 33.4 - 50.7 - 78.9 1212.8 - - - - - Shikra  [88] Vicuna-13B  [84] - 73.9 117.5 77.4 - - - - - - - 58.8 - - - 54.7 InternVL-Chat*  [63] Vicuna-13B  [84] 126.2 92.2 146.2 81.2 66.6 58.5 - 61.5 - 87.6 1586.4 - - - - - Qwen-VL  [89] Qwen-7B  [89] 121.4 85.8 - 78.8 59.3 35.2 67.1 63.8 - - - 38.2 - 56.3 - 59.4 Qwen-VL-Chat  [90] Qwen-7B  [89] 120.2 81.0 - 78.2 57.5 38.9 68.2 61.5 - - 1487.5 60.6 - 58.2 - - TinyGPT-V  [91] Phi2-2.7B  [92] - - - - 33.6 33.4 - - - - - - - - - - LLAVA-Phi  [93] Phi2-2.7B  [92] - - - 71.4 - 35.9 68.4 48.6 - 85.0 1335.1 59.8 - - 28.9 - BLIP-2  [94] FLAN-T5  [56] 103.9 71.6 - 41.0 41.0 19.6 61.0 42.5 - 85.3 1293.8 - 38.1 46.4 22.4 -
*: InternVL-Chat with IViT-6B as the visual encoder, QLLaMA as the glue layer and 13B trained parameters
Bold: The best results
Underlined: The second best results
@@ 3 Key technologies for building domain-specific foundation models
This section delves into the technical pathways toward customizing domain-specific FMs. We will explain how to flexibly select and combine the appropriate modules from five key components — MEs, IPs, BCs, OPs, and MDs based on the specific requirements of different domains. Additionally, we will analyze concrete cases to help readers better understand and apply the methodologies discussed in this section.
We can categorize the customization of domain-specific FMs into three levels, ranging from low to high level of customization (i.e., from high to low reliance on general-purpose FMs or pre-trained modules):
@@
    Domain-specific enhancement based on general-purpose FMs.
@@
    Customization of the FM based on pre-trained modules.
@@
    Construction of the FM without pre-trained modules.
Table 4 summarizes the characteristics of these three domain-specific FM customization methods.
Table 4: Customization methods of domain-specific FMs
@@
Based on general-purpose FMs Low, only the input method of domain knowledge is customized Low Low Low
@@
3.1 Domain-specific enhancement based on general-purpose foundation model
General-purpose FMs offer comprehensive capabilities, making them suitable for various task scenarios. When such a general-purpose FM can fully handle the required data modalities, any modifications to its underlying architecture for model developers are rendered unnecessary. Instead, they can focus on implementing domain-specific enhancements.
Depending on whether the domain-specific enhancement requires altering the parameters of the FM, we can further divide this into plug-and-play domain-specific enhancement and fine-tuning-based domain-specific enhancement. Table 5 categorizes and summarizes the methods of domain-specific enhancement based on a full-architecture general-purpose FM.
Table 5: Specific-domain enhancement with the entire general-purpose FMs
@@ 3.1.1 Plug-and-play domain-specific enhancement
The generality, generalization capabilities, and reasoning abilities of general-purpose FMs enable them to serve as the foundation for domain-specific models. To achieve plug-and-play domain enhancement without modifying the parameters of the FM, two approaches can be utilized: leveraging existing knowledge or embedding new knowledge. The first approach aims to leverage domain knowledge already stored within the general-purpose FM, as illustrated in Figure 6 (a). The second approach, embedding new knowledge, involves equipping the FM with the ability to handle domain tasks by introducing domain-specific knowledge. This can be divided into two methods: embedding knowledge through prompts and embedding knowledge via an external knowledge base. These methods are depicted in Figure 6 (b) and (c) respectively. The following sections will provide a detailed explanation of these techniques.
@@
    Invoking existing knowledge for domain enhancement: A general-purpose FM may have already enclosed domain knowledge during the training process. Prompt tuning improves prompts to better invoke the model’s inherent domain knowledge, where ”tuning” refers to the optimization of prompts. Specifically, it involves inserting a carefully crafted prompt into the input data as context to improve the generated output. These carefully crafted prompts can be natural language descriptions, examples, rules, or other text or embedding vectors that guide the model to understand the task requirements. The model will consider these carefully crafted prompts to produce task-relevant results when generating outputs. Prompt tuning can be categorized into hard prompts and soft prompts:
@@
      Hard prompts: Hard prompt methods are common techniques in natural language processing (NLP). They use interpretable and reusable handcrafted words and tokens to guide the language model’s output. Hard prompts are typically manually designed and tailored for specific tasks, making them difficult to modify. PET (Pattern Exploiting Training) [95] is a classic hard prompt learning method that models questions as cloze tasks and optimizes the final output words. This method trains the model on a small amount of supervised data and performs ensembling predictions on unsupervised data to guide the model.
@@
      Soft prompts: Designing hard prompts requires experimental exploration and expertise, and manually designed prompts may not align well with the model’s data processing methods. To simplify this process and enhance the flexibility of prompt tuning, researchers proposed soft prompt-based tuning methods. Prefix Tuning [96] is a form of soft prompt tuning that adapts to specific downstream tasks by adding learnable prefix vectors (soft prompts) to the beginning of the input sequence. As part of the input, these prefix vectors guide the model’s output to meet task requirements. The advantage of prefix tuning is that it only updates these prefix vectors rather than the model’s parameters, significantly reducing computational and storage resource demands while retaining the rich knowledge learned by the pre-trained model. Building on prefix tuning, researchers introduced the P-tuning method [97]. P-tuning replaces fixed or manually designed words and tokens with learnable soft prompts. Its core idea is to treat prompts as part of the model that can be learned, allowing the model to learn not only how to respond to given tasks but also how to generate the best prompts. These soft prompts are typically a series of embedding vectors processed alongside the actual text input. Through end-to-end training, the model automatically learns to adjust these embedding vectors for better task performance. P-tuning combines the parameter efficiency of prefix tuning with the flexibility of traditional hard prompt tuning. Soft prompts give the model greater freedom to generate answers, potentially producing more diverse outputs and increasing the risk of generating inaccurate or irrelevant responses.
@@
    Knowledge embedding for domain enhancement: When the existing knowledge of a general-purpose FM is insufficient to solve domain tasks, introducing new knowledge by embedding additional background information can achieve higher-quality output. This method is known as knowledge embedding for domain enhancement.
@@
      • Resource consumption of the attention mechanism: The attention mechanism is the core of Transformer models, allowing the model to compute attention weights for each element in the sequence. However, as the sequence length increases, this mechanism’s computational complexity and memory requirements grow quadratically, leading to significant resource consumption.
      • Long-distance dependency issue: When handling long sequences, the Transformer needs to span a large number of input tokens, which often results in problems such as gradient vanishing or exploding. This makes it difficult for the model to capture dependencies between elements far apart in the sequence.
      To address the above issues, Lossless Long Text technology has emerged. It aims to enhance the model’s ability to process long texts that exceed its input length limitations, allowing users to input a large amount of domain knowledge directly through prompts into the large language model as contextual information for domain enhancement. Lossless Long Text technology expands the long-text input capability of large language models in two directions: extrapolation and interpolation:
@@
      Figure 7: An example of using LLM to answer multiple-choice communication questions. RAG improves the accuracy of answers by retrieving additional knowledge for the LLM from a knowledge base based on the relevance to the question text. The data used in the example is from [114].
@@
      Knowledge embedding by external knowledge base: In practical applications, users may be unable to provide sufficient domain knowledge to enhance a general-purpose FM. To address this issue, model deployers can augment the general-purpose FM with a dedicated domain knowledge base. This method allows the general-purpose FM to reference this external knowledge base when generating answers or performing tasks, thereby obtaining the necessary domain information and context to provide more accurate and targeted responses or solutions. Retrieval-Augmented Generation (RAG) [100, 1, 115] technology was developed to achieve this purpose. RAG technology aims to enhance the language model’s generation capabilities by leveraging an external document base without retraining the model. It is particularly suitable for tasks requiring a customizable dynamic knowledge base, such as question-answering, text summarization, and fact-checking. The core of RAG technology is the integration of a retrieval component that can quickly find information relevant to the current task in a large document database during the generation process. Once the relevant documents are retrieved, this information is used as additional contextual information to aid the generation process. The advantage of RAG technology is that it combines the generation capabilities of large language models with the knowledge provided by external retrieval systems without requiring domain knowledge themselves. Furthermore, because the external knowledge base can be replaced as needed, RAG technology offers high flexibility and adaptability. To have an intuitive understanding of RAG, we present an example in Figure 7.
@@
      In RAG, the two core challenges are retrieving useful information from the database and organizing that information to augment the generation of FMs. For the retrieval task, the earliest and most naive approach uses sparse retrieval that directly matches data based on raw data like BM25 [116]. Inspired by the field of information retrieval, dense retrieval like DPR [117] has been proposed, which projects raw data into a high-dimensional space, potentially helping to capture semantic similarity better. However, a significant challenge is that even if the retrieved data is highly related to the original input in the semantic space, we cannot guarantee that it will help the model generation due to issues like polysemy. To address this problem, some researchers have introduced other technologies like knowledge graphs to aid the retrieval process [101]. [118] showed that adding noise can help enhance the model performance compared to traditional dense retrieval methods. Additionally, the organization of the retrieved information can directly influence the output quality. For example, providing a ranking of the retrieved data [102] can improve model performance.
@@
While the aforementioned technologies were initially proposed to enhance large language models in specific domains, their applications are not limited to language models. With the development of FM domains, these technologies are expected to be extended to FMs of other modalities.
@@ 3.1.2 Domain-specific enhancement based on fine-tuning
When plug-and-play domain enhancement techniques are difficult to implement or require embedding too much domain knowledge into the general-purpose FM, or when deep modifications to the general-purpose FM are necessary, we can turn to the strategy of domain enhancement based on fine-tuning. This strategy aims to customize the required domain-specific FM while preserving the pre-trained knowledge of the general-purpose FM as much as possible by specific domain enhancement [119].
Fine-tuning techniques can be divided into three main types: adapter-based fine-tuning, low-rank matrix decomposition-based fine-tuning, and full-parameter fine-tuning. Figure 8 (a), (b), and (c) illustrate these three technical pathways, respectively. In the following sections, this paper will elaborate on these techniques, sorted from low to high resource requirements and the complexity needed for fine-tuning.
@@
    Adapter-based fine-tuning: Adapter-based fine-tuning [103] is a method that inserts small trainable adapter modules into pre-trained models, aiming to efficiently adapt the model to specific downstream tasks. During fine-tuning, only the parameters of adapter modules are updated, while the original parameters of the pre-trained model remain unchanged, reducing computational resources and storage requirements while retaining the rich knowledge learned by the model during pre-training. AdapterFusion [104] is an extension of adapter-based fine-tuning that allows the model to learn multiple tasks or adapt to various data distributions simultaneously by fusing multiple adapter modules, with each adapter module focusing on capturing task-specific features. Infused Adapter by Inhibiting and Amplifying Inner Activations (IA3) [105] scales the activation layers by injecting learned vectors into the attention and feed-forward modules of the Transformer architecture. Since these learned vectors are only trainable parameters during fine-tuning, IA3 significantly reduces the number of trainable parameters compared to traditional adapter-based fine-tuning, thereby reducing training costs and improving training efficiency. Additionally, IA3 does not introduce inference latency because its adapter weights can be merged with the FM while maintaining the flexibility and adaptability of the model, allowing customized fine-tuning for different tasks and datasets. [106] proposes a method called model reprogramming that converts both the input and output domains of an FM by adding two adapter layers, enabling more flexible adaptation. Unlike inserting adapters into each layer in the FM, model reprogramming serves as a more direct way in which only the input and output adapters are inserted upstream and downstream of the FM. During reprogramming, the pre-trained FM is frozen, which means the core of the model still works in the original domain. Such a process reduces the cost of training resources and enables the simultaneous adaptation of a single pre-trained FM to multiple domains. However, because of less trainable adapters, such a method depends more on the similarity between the source domain of the pre-trained FM and the target domain. For example, time-series data can be considered similar to text data since they all persist in a series form [17]. To better illustrate the ability of model reprogramming, we present an example in Figure 9.
    Figure 9: An example of model reprogramming. The input time-series data is embedded by a patch reprogram module, which patches the data and then projects each patch embedding into the text space. Besides the time-series part, text information of the dataset is also inputted for LLM to better understand the prediction task. The data used in the example is from [17].
@@
    Low-rank matrix decomposition based fine-tuning: Fine-tuning based on low-rank matrix decomposition reduces the number of parameters that need to be updated by decomposing the weight matrices in the pre-trained model into the product of low-rank matrices. Low-rank matrix decomposition can capture the most important information in the weight matrices while keeping the original pre-trained parameters unchanged during fine-tuning, updating only low-rank decomposition matrices, thus reducing the computational and storage requirements during fine-tuning. This method improves fine-tuning efficiency while maintaining or approaching the performance of full-parameter fine-tuning. In low-rank matrix decomposition-based fine-tuning methods, two low-rank matrices, $A$ and $B$, share the same rank $r$, are introduced to construct an update parameter matrix. The low-rank property of both $A$ and $B$ enables this process to be computationally efficient while still enabling the model to adapt effectively to new tasks.
@@
      Low-Rank Adaptation (LoRA): Low-Rank Adaptation (LoRA) [107] achieves efficient fine-tuning performance by introducing low-rank adaptation matrices to the pre-trained model, rather than updating the original weight matrix directly. LoRA modifies the model by adding a low-rank update to the original weights, as described by the equation:
@@
      where $W_{\text{old}}$ is the original weight matrix, $\Delta W$ is the low-rank update matrix, and $+$ is the point-wise addition of two matrices with the same shape. The key idea is that $\Delta W$ is constructed as the product of two low-rank matrices $A$ and $B$ , such that:
      $\Delta W=AB\text{.}$ (11)
      Here, $A$ and $B$ are low-rank matrices with ranks typically much smaller than that of $W_{\text{old}}$. During fine-tuning, only the matrices $A$ and $B$ are updated, while $W_{\text{old}}$ remains fixed. These low-rank matrices are learned to adapt the model efficiently while keeping the overall number of trainable parameters minimal.
@@
      Researchers have also found that LoRA’s continuous pre-training performance is unsatisfactory on some large-scale datasets. Thus, the Layerwise Importance Sampled AdamW (LISA) [121] strategy is proposed, where the weight norm distribution of different layers exhibits uncommon skewness. LISA adopts an importance sampling strategy by randomly activating different layers in the FM for optimization. Specifically, LISA consistently updates the bottom embedding layers and the top linear head while randomly updating a small number of intermediate self-attention layers. With memory consumption comparable to LoRA, this method outperforms LoRA and even full-parameter fine-tuning in various downstream fine-tuning tasks.
@@
      $W_{\text{new}}=W_{\text{old}}+\Delta W\text{,}$ (12)
      where $W_{\text{old}}$ is the original weight matrix and $\Delta W$ is the update matrix approximated by a low-rank matrix. The update matrix $\Delta W$ can be further decomposed into the Hadamard product of two low-rank matrices:
      $\Delta W=A\odot B\text{.}$ (13)
      Here, $A$ and $B$ are two low-rank matrices that adjust elements of the original weight matrix by learning key information extracted from the input data, and $\odot$ is the Hadamard product. By updating only parameters in $A$ and $B$, LoHa achieves efficient model fine-tuning while maintaining adaptability to new tasks.
@@
      Low-Rank Kronecker Product (LoKr): Low-Rank Kronecker Product (LoKr) [109] is another parameter-efficient fine-tuning method that emerged after LoHa. LoKr utilizes the properties of the Kronecker product to expand the dimensions of the weight matrix while keeping the increase in the number of parameters within a manageable range. The Kronecker product allows the model to learn complex interactions across different dimensions, particularly useful for capturing high-order relationships in the input data. The updating process of LoKr can be represented as: